OpenStudy (wolfe8):

Need help guessing the particular solution form for y'''+y''-2y=te^(-t)+1

OpenStudy (anonymous):

First the homogenous solution: You have the characteristic equation, \[r^3+r^2-2=0\] which has roots \(r=1, -1\pm i\). The complex roots' solutions are of no consequence to your guess, since they contain factors of sine/cosine, and \(r=1\) contributes a \(e^t\), which also won't affect your guess solution. I'd try \(y_p=Ate^{-t}+Be^{-t}+Ct+D\). \[y'=(A-B)e^{-t}-Ate^{-t}+C\\ y''=(B-2A)e^{-t}+Ate^{-t}\\ y'''=(3A-B)e^{-t}-Ate^{-t}\] \[\left[(3A-B)e^{-t}-Ate^{-t}\right]+\left[(B-2A)e^{-t}+Ate^{-t}\right]\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~-2\left[Ate^{-t}+Be^{-t}+Ct+D\right]=te^{-t}+1\\ (A-2B)e^{-t}-2Ate^{-t}-2Ct-2D=te^{-t}+1\] Matching up coefficients, you have \[\begin{cases}A-2B=0\\ -2A=1\\ -2C=0\\ -2D=1\end{cases}\] I've checked the answer, it all works out. Note that you can disregard the \(Ct\) part of the guess, I only used it to be safe.

OpenStudy (wolfe8):

Thanks a lot man!